OpenStudy (anonymous):

evaluate the integral: ∫(x+4)/((x^2)+2x+5)dx

OpenStudy (anonymous):

i got: \[\frac{ 1 }{ 6 }\ln (x+3)+\frac{ 6 }{ 5 }\ln (x-2)+c\] ??????